OpenStudy (anonymous):

Which is a solution of the equation? 2x – y = 5

OpenStudy (anonymous):

A. (2, –1) B. (3, –1) C. (1, 2) D. 7

jimthompson5910 (jim_thompson5910):

Let's check to see if choice C is a solution or not

jimthompson5910 (jim_thompson5910):

To do this, we simply plug x = 1 and y = 2 into 2x – y = 5. This is because (1,2) is of the form (x,y) 2x – y = 5 2(1) – 2 = 5 ... plug in x = 1, y = 2 2 - 2 = 5 0 = 5 ... this is FALSE, so (1,2) is NOT a solution to 2x – y = 5

jimthompson5910 (jim_thompson5910):

Using these steps, is (3, –1) a solution?

OpenStudy (anonymous):

thanks dude but thats confusing

jimthompson5910 (jim_thompson5910):

which part doesn't make sense?

OpenStudy (anonymous):

mostly all of it

jimthompson5910 (jim_thompson5910):

do you see how I got x = 1, y = 2 from the point (1,2) ?

OpenStudy (anonymous):

never mind I understand it had to clean my glasses

OpenStudy (anonymous):

I* not it